OpenStudy (anonymous):

Luther purchased a motorcycle for $11,592. It depreciates about 7.2% each year. What is the value of the motorcycle after six years

OpenStudy (anonymous):

1st you nee to find out how much money it will lose in value each year. 11,592*0.072=834.624 2nd multiply what you lose a year but total of year you losing money for. 834.624*6=5,007.744 3rd is to subtract to get the total amount that motorcycle would now be worth. 11,592-5,007.744=6,584.256 so after 6years the motorcycle would be worth $6,584.26

OpenStudy (anonymous):

aww thnx for walking me through the steps i appreciate it

OpenStudy (anonymous):

no problem. :)